A newly discovered particle is found to have a quark composition of anti-bottom, up. What is the charge of this particle, in Coulombs?


1
Expert's answer
2021-06-23T12:41:37-0400

Gives

Particle found quark composition

="(\overline{B} u)"

Charge of anti bottom quark

"\overline{B}=\frac{1}{3}e"

Charge up quark

"u=\frac{2}{3}e"

quark composition net charge "=\frac{1}{3}e+\frac{2}{3}e=e"

Charge of partical "(\overline{B} u)=e=1.6\times10^{-19}C"

"q_{(\overline{B} u)}=1.6\times10^{-19}C"